OpenStudy (anonymous):

Anyone could HELP me with this please.... How to find the volume of the solid S from: a right circular cone with height h and base radius r

OpenStudy (anonymous):

What I try is.. as it was found that \[r = \sqrt{c ^{2}-h ^{2}}\] so the cross sectional area is A(h)= pi r^2 = pi (c^2-h^2) from the drawing.. a=0, b=h so the volume is \[\int\limits_{0}^{h}A(h) dh = \int\limits_{0}^{h} pi*(c ^{2}-h ^{2}) dh = 1/3 pi * r ^{2} h\] but this wasn't quite right.

OpenStudy (anonymous):

\[\[\int\limits\limits_{0}^{h}A(h) dh = \int\limits\limits_{0}^{h} pi*(c ^{2}-h ^{2}) dh = 1/3 pi * r ^{2} h\]\]

OpenStudy (anonymous):

i have learned a formula as 1/3 pi r^2 h.... and i was able to solve problems with this.

OpenStudy (anonymous):

@reeba @Rohangrr thanks everyone ;)